OpenStudy (anonymous):

what are the x-intercepts and minimum point(s) for the equation 3x^2-13x-10

OpenStudy (amistre64):

minimum is when x = 13/6

OpenStudy (anonymous):

Set the equation equal to zero to find the x intercepts. (When y=0 the graph is on the x axis). Using the quadratic equation you end up getting 5 and -2/3. (You can verify this yourself). To find the minimum you can use calculus or algebra, but seeing as amistre64 answered this, he can explain it.

OpenStudy (amistre64):

we can all explain it ;)

OpenStudy (anonymous):

I can't do it with algebra :). Calculus drains all basic algebra from the brain. Haha.

OpenStudy (amistre64):

dont it tho lol

OpenStudy (amistre64):

the max/min of teh graph of a quadratic equation is where the graph curves back on itself and backs a peak or valley; that point is valled the vertex of the parabola

OpenStudy (amistre64):

teh definition of a parabola also helps to define the vertex as the point in line between the focus and the directix such that the line from the focus to the directix is perpendicular it passes thru the vertex

OpenStudy (amistre64):

the distance between all points of a parabola are equal distance to the focus and a perpendicular line to the directex; the shortest distance shich results is called the vertex

OpenStudy (amistre64):

and yada yada yada ...lol

OpenStudy (radar):

I don't know what method was used to get the 13/6 but one method is to\[d/dx=6x-13\]

OpenStudy (amistre64):

-b/2a ;)

OpenStudy (radar):

6x-13=0 x=13/6

OpenStudy (amistre64):

the determinate of the quadratic formula is the x axis of symmetry if the parabola is oriented such that it is a function of x

OpenStudy (amistre64):

(-b/2a, f(-b/2a)) are the coordinates of teh vertex....

OpenStudy (amistre64):

there is also a method called the square root method that might be applicable

OpenStudy (anonymous):

Refer to the plot attachment.

OpenStudy (anonymous):

x=-.66666667 & x=5 the min (vertex) s x=2.1, y=-24.07